Solve the equation:
(1/a+3)-(a/a^2-9)=(2/3-a)

Answer by ewatrrr(24785)

 
Hi,
(1/a+3)-(a/a^2-9)=(2/3-a)
1/(a+3)- a/(a+3)(a-3)= -2/(a-3) |Note Parenthesis used to show terms in denominator
|Multiplying thru by (a+3)(a-3) so as all denominators = 1
(a-3) - a + 2(a+3) = 0
2a +3 = 0
2a = -3
a = -3/2